The next dividend payment by Hoffman, Inc., will be $3.05 per share. The dividends are anticipated to maintain a growth rate of 6 percent forever. Assume the stock currently sells for $49.70 per share. a. What is the dividend yield? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) b. What is the expected capital gains yield? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a) Calculation of Dividend yield:
Dividend yield= dividend/current price*100
Dividend yield= 3.05/49.70*100= 6.14%
Dividend yield= 6.14%

b) calculation of capital gain yield:
Price after year 1(P1)= current price*(1+growth)= 49.70*(1+0.06)= 52.682

Capital gain yield= (P1-P0)/P0*100
Capital gain yield= (52.682-49.70)/49.70*100= 6%

Capital gain yield= 6%